lj2005:

Alina is able to swim at a rate of about 2.5 miles per hour. Talia is able to swim at a rate of 1.8 meters per second. Which statement comparing the two swimmers is accurate?

jhonyy9:

any idea ? may be possibile a swimmer swim miles on a concurs ?

lj2005:

not at all

jhonyy9:

yes sure so then how you think what may be the correct answer ?

jhonyy9:

so then not is miles per hour so what remained there ?

Mercury:

answer choices: 1 mile 1,609 meters A) Alina swims about 0.7 miles per hour faster than Talia. B) Talia swims about 0.7 miles per hour faster than Alina. C) Alina swims about 1.5 miles per hour faster than Talia. D) Talia swims about 1.5 miles per hour faster than Alina.

Mercury:

since the answer choices are in miles/hr, it would be wise to convert Talia's speed from m/s to miles/hr setting up the dimensional analysis \[\frac{ 1.8~meters }{ \sec }\frac{ 1~mile}{ 1609~meters}\frac{ 3600~\sec }{ 1~hr}\]